The length of a rectangle is 1 inches less than three times the width. If the length is increased by 6 inches and the width is increased by 5 inches, then the length will be twice the width. Find the dimensions of the rectangle.
Answer by ptaylor(2198)

Let W=width
Then length, L=3W-1------------------eq1
we are also told that:
L+6=2(W+5)or
L+6=2W+10 subtract 6 from each side
L+6-6=2W+10-6 collect like terms
L=2W+4------------------------------eq2
Substitute L=2W+4 from eq2 into eq1:
2W+4=3W-1 subtract 2W from and add 1 to both sides
2W-2W+4+1=3W-2W-1+1 collect like terms
5=W or
W=5 in ------------------------ans for W
substitute W=5 in into eq1
L=3*5-1 or
L=15-1=14 in-----------------------ans for L
CK
14=3*5-1
14=14
and
14+6=2(5+5)
20=20
Hope this helps---ptaylor